OpenStudy (anonymous):

Please help! I don't get any of this! :( 1. What is the slope of in the graph below? A. -2 B. -1 C. 1 D. 2 2. Find the slope of a line that passes through (-2,-3) and (1,1). A. 1/3 B. 1 C. 2 D. 4/3

OpenStudy (anonymous):

where is the graph

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

There

OpenStudy (anonymous):

For the first one, the slope is -2, or -2/1. As you can see, the rise is -2, and the run is 1. Rise always goes over the run. If the slope has the number 1 underneath the rise, you can simply just say that the slope is the rise, like the -2. So, the answer is B. -2. Does that clear things up?

OpenStudy (anonymous):

Yes thank you :)

OpenStudy (anonymous):

You're welcome! (:

sammixboo (sammixboo):

Slope = \(\frac{rise}{run}\) The rise is already given, which is -2, and the run is 1, so we replace -2 with tise and 1 with the run \(\frac{-2}{1}\) = -2

sammixboo (sammixboo):

I am lagging so much

OpenStudy (anonymous):

3-1 / 0 - 1 = -2 ( the first question) y2-y1/x2-x1

OpenStudy (anonymous):

-3 - 1 / -2 - 1 ) -4 / -3 = 4/3 so the answer is D

OpenStudy (anonymous):

so 1 is D?

OpenStudy (anonymous):

For the second one, use the slope formula. \[\frac{ y2-y1 }{ x2 -x1}\] Plug in the two points. (-2, -3) (1, 1) 1 is y2, and -3 is y1. 1 is x2, and -2 is x1. \[\frac{ 1 - (-3) }{ 1 - (-2) } = \frac{ 1 + 3 }{ 1 + 2 } = \frac{ 4 }{ 3 }\] So, the answer is D. 4/3. It's really simple when you memorize the formula, huh? (:

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (anonymous):

Summary: (I said B. -2 for the first one, but -2 is A, not B haha, sorry.) So, the first one is A. -2, and the second one is D. 4/3

OpenStudy (anonymous):

I'll try. Close this first and ask the questions in a new post. (:

OpenStudy (anonymous):

Okay will do